what amount of sucrose will contain same number of atom as the number of ion in 2.84 mg of Na2SO4

Grade:11

1 Answers

Vikas TU
4 years ago
Dear student 
Mol. wt. of sucrose = 342 gm
Mol .wt of Na2SO4 = 142 gm
Moles = 2.84 / 142 = 0.02
No. of atoms = 0.02 NA
According to question
0.02 NA = x / 342 NA
0.02 = x / 342
x = (2 / 100)*342
x = 684 / 100
x = 6.84 gm
Hope this helps 
Good luck